OpenStudy (anonymous):

Multiply : 3x+13/8m * m^2/2x+8 3m/16 , m Not= 0 3m/16 , m not = 0 and xNot= -4 m(x+3)/8 , m Not=0 m(x+3)/8 , m Not=0 and xNot=-4

OpenStudy (anonymous):

LMAO

OpenStudy (anonymous):

Will you qctually help me please ?

OpenStudy (anonymous):

LOL

OpenStudy (across):

I will help, tutor, friend and raise you, heck, I'll even write you a book, if you first show me what you've tried.

OpenStudy (anonymous):

I don't get it though ...

OpenStudy (across):

Okay! "Multiply : 3x+13/8m * m^2/2x+8" First, this looks much better when you format it: Multiply : \(\left(3x+\frac{13}{8}m\right)\cdot\left(\frac{m^2}{2m}+8\right)\) Is this what you're trying to solve?

OpenStudy (anonymous):

No ... (3x+13)/(8m) * (m^2)/(2x+8) Sorry about that

OpenStudy (anonymous):

No ... (3x+13)/(8m) * (m^2)/(2x+8) Sorry about that

OpenStudy (across):

\[ \frac{3x+13}{8m}\cdot\frac{m^2}{2x+8}\text{ This is the problem.}\\ \frac{3x+13}{8}\cdot\frac{m}{2x+8}\text{ The }m\text{'s cancel.}\\ \frac{3xm+13m}{16x+64}\text{ Multiply top with top and bottom with bottom.}\\ \]

OpenStudy (anonymous):

The 3rd option is the answer ? ?

OpenStudy (anonymous):

u can see clearly here http://www.wolframalpha.com/input/?i=3x%2B13%2F8m+*+m%5E2%2F2x%2B8

OpenStudy (anonymous):

u check it

OpenStudy (anonymous):

Yeah . that wasnt the right problem ... I pot the right problem in a comment above